Obtain the quartile of the data 33, 25, 42, 25, 31, 37, 46, 29, 38 


1
Expert's answer
2021-09-01T10:51:01-0400

"\text{Rearranging the data we have: }25,25,29,31,33,37,38,42,46 \text{ number of data (n) = 9}\\
\text{first quartile ($Q_1$) = $\frac{n + 1}{4}$th term $= \frac{10}{4} = 2.5$th term to get the 2.5th term we add the 2nd and the 3rd term and divide by 2.} \\
\implies Q_1 = \frac{25 + 29}{2} = 27\\
\text{second quartile ($Q_2$) = $\frac{n + 1}{2}$th term $= \frac{10}{2} = 5$ th term = 33 } \\
\text{third quartile ($Q_3$) = $\frac{3(n + 1)}{4}$th term $= \frac{30}{4} =7.5$th term to get the 7.5th term we add the 7th and the 8th term and divide by 2.} \\
\implies Q_3 = \frac{38 + 42}{2} = 40"